OpenStudy (anonymous):

what is the simplest way to solve this equation? 55^3 + 25^3 - 35^3 = 155000

Parth (parthkohli):

A calculator!

OpenStudy (anonymous):

55=5*11 25=5*5 35=5*7 so take out the common factor of 3 first term 5^3

OpenStudy (anonymous):

It is always a calculator...

OpenStudy (anonymous):

@geedee do some reply...

mathslover (mathslover):

\[\large{55^3 + 25^3 -35^3}\] \[\large{(5*11)^3+(5*5)^3-(5*7)^3}\] \[\large{5^3(11^3+5^3-7^3)}\] \[\large{125(1331+125-343) = 125(1113) = 139125}\]

Parth (parthkohli):

\[(5x)^3 + (5y)^3 - (5z)^3 \]\[125x^3 + 125y^3 - 125z^3 \]\[125(x^3 + y^3 - z^3) \]Does that help?

Parth (parthkohli):

Always take variables from algebra when you are confused about the constants. :)

OpenStudy (anonymous):

@ParthKohli be yourself...

Parth (parthkohli):

Erm?

mathslover (mathslover):

or use : \[\large{a^3+b^3+c^3=(a+b+c)(a^2+b^2+c^2-ab-bc-ca)+3abc}\] \[{[(55+25-35)(55^2 +25^2 +(-35)^2-(55)(25)-(25)(-35)-(-35)(55)]+3(55)(25)(-35)}\]

OpenStudy (anonymous):

I am talking about your pic and your first comment too where you have written : "does that help??"

OpenStudy (anonymous):

But that is not the simplest way @mathslover

Parth (parthkohli):

I'm not perfectly sure what you are talking about... :I

mathslover (mathslover):

either that is not a simple way... though I posted so that the user is satisfied with each and every step or process

mathslover (mathslover):

Must say that the process I mentioned earlier is the easiest way after using calculator

OpenStudy (anonymous):

Nevermind but Be Yourself @ParthKohli

Parth (parthkohli):

The simplest way is to use a calculator :) *sigh*

OpenStudy (anonymous):

If you are allowed to use calcualator then yes calculator is the simplest...

Parth (parthkohli):

Everyone is allowed to use technology: that is what we made it for. :)

OpenStudy (anonymous):

i think formula that mathlover used is simple and easy to remember.

mathslover (mathslover):

that's nice to hear but @geedee that involves too much calculation

OpenStudy (anonymous):

Today, my teacher gave me lecture on Technology.. And He said, technology has affected our Human Values and he is pretty correct in that..

OpenStudy (anonymous):

He said, Graph of Technology is increasing day by day but Human values are decreasing at the same rate at which Technology is increasing...

Parth (parthkohli):

@waterineyes by the way, please watch this. http://www.youtube.com/watch?v=60OVlfAUPJg

OpenStudy (anonymous):

There is no full access in this computer lab of my college...

Parth (parthkohli):

@geedee please watch that talk :)

OpenStudy (anonymous):

Ha ha ha... Thanks geedee.. By the way same teacher has taught you too geedee I know that...

OpenStudy (anonymous):

And I think in your lab too full access is not there..

OpenStudy (anonymous):

Ha ha ha...